# Equal contributions

### v2

@@ III-D Other approaches: Parallel sums with OpenMP
OpenMP, a directive-based API for shared memory and accelerator parallel programming, supports data reduction to perform parallel calculations that are portable across architectures. The OpenMP programming model uses reduction scoping clauses to specify the regions of code where the reduction takes place. These regions can include parallel fork-join, tasks, sections, SIMD, and scope constructs [23], which can be executed on a host or a device such as a GPU. The OpenMP specification does not specify the location and ordering in which the values are combined, thus, bitwise determinism is not guaranteed. As a result, implementing a deterministic parallel reduction in OpenMP will require using constructs that can enforce ordering when reducing private and shared variables. An ordered construct associated with a block of can define a structured block within a loop, SIMD, or loop SIMD region. This construct enforces sequential execution for a specified region of code according to the loop iterations, while enabling parallel execution for code outside the region. When the thread handling the first iteration of the loop reaches the ordered construct, any other thread that encounters the ordered construct in its iteration will wait at the start of the ordered region until all ordered regions from previous iterations have been executed. To order the entire loop, an ordered clause can be used.
@@
Listing 3: An ordered parallel reduction implementation OpenMP
@@
Listing 4: An ordered parallel reduction implementation OpenMP using
@@
Listing 3 shows the use of a block-ordered directive on a reduction. A similar effect can be achieved while using an ordered clause. Listing 4 shows how to use an ordered clause inside a target region that executes on a device or GPU. When comparing the results of listing 3 with a reduction without the ordered directive, using gcc 12.2.0 on CPU, we get the results shown in Table III; the ordered version produces deterministic results.
TABLE III: Results of the normal and ordered reductions
@@
IV Non-determinism in PyTorch
@@
Apart from the choice of kernel at runtime and atomic operations, there are several other ways in which PyTorch may induce variability. These include having an unset random number generator seed, having unset CUDA environment variables, uninitialized GPU memory and communication between devices. To focus exclusively on variability emergent from the first two sources, we use PyTorch environment variables and a single GPU to remove those other sources of variability. We control whether or not non-deterministic kernels are available through an environment variable [24]. We note that this documentation is not necessarily accurate across all systems and software versions, as we received a runtime error when trying to obtain a deterministic result for scatter_reduce.
@@
We performed experiments in two ways, depending on whether or not a deterministic kernel exists. If there is a deterministic option, then the output multidimensional array $\mathbf{A}$ is fixed by that implementation. We compare it with the tensors $B_{i}$ with $i$ labelling $N$ runs of the non-deterministic implementation. If there is no non-deterministic kernel, we choose the first non-deterministic invocation as a reference, $\mathbf{A}=\mathbf{B}_{0}$. For each experiment we also present results for measurements of kernel runtime for non-deterministic implementations on the GPU, and deterministic implementations on the GPU and LPU architectures, to expose the performance costs of using deterministic operations. Runtime measurements are only for the execution time of the operation, excluding data transfers to/from the device. For the GPU we make many measurements to report the average and standard deviation of the execution time. On the LPU architecture the runtime for the operation is reported as a fixed number since the cycle-by-cycle execution is determined ahead of time [25]. We performed a sweep over operation hyperparameters with $N=1,000$ runs on an H100 and observed only a handful of operations specified in the PyTorch documentation[24] to exhibit non-determinism. We list these operations in Table IV alongside minimum and maximum $V_{\text{ermv}}$. We refer the reader to our GitHub [22] for in-depth details on the full hyperparameter sweep, where, for example, we consider kernel size, stride and padding when testing convolution operations.
@@
Our test results indicated that the main factor in ability to observe output variability in these kernels is the size of the input tensor, which gives more chances to observe the switching order of atomic operations at runtime. For operations which perform reductions on an input tensor, we found that the ratio of the output dimension to input dimension is another key factor. We explore this further, restricting our analysis to scatter_reduce and index_add, which have such reductions.
@@
TABLE IV: Max and min variability for non-deterministic pytorch operations over all hyperparameters tested.
Operation min$(V_{\text{ermv}})\times 10^{-7}$ max$(V_{\text{ermv}})\times 10^{-6}$
@@ IV-A Case studies of PyTorch kernels
The scatter_reduce operation updates an output array $Y$ by applying a reduction operation on values from the input array $X$, according to indices specified in an index array $I$, for example, when reducing a one dimensional array with summation reduction, $Y[i]=\sum_{j}\left(\left\{X[j]\mid j=I[i]\right\}\right)$. This can be generalized to arrays with arbitrary dimensions and different reduction functions. The index_add operation updates the an output array $Y$ by adding values from an input array $X$ according to the indices specified in an index array $I$. As an example, for two dimensional input and output arrays with a summation reduction over the first dimension, the elements of $Y$ are $Y[i,j]=Y[i,j]+X[k,j],\text{ for}\ i=I[k],\forall k$. This can similarly be generalized to arbitrary dimensions and different reductions other than addition. Both of these operations reduce the size of a input tensor, so we define the reduction ratio $R$ as the ratio of the output dimension size to source dimension size. When $R=1$, the arrays have the same size, and smaller values of $R$ correspond to larger reductions of the source array down to at most a singular value, along the axis of reduction. For the index tensors of both operations, we use random integers drawn from a uniform distribution to choose arbitrary values from the source tensor. The motivation for this is to mimic an arbitrary graph structure, where the reduction is happening over nodes that share an edge, although in this case we do not a priori expect any specific structure or hierarchy to be emergent.
Figure 3: Heatmaps of unique outputs per run/iteration for 1,000 runs of the non-deterministic implementation of scatter_reduce  (left) and index_add  (right) for different reduction ratios and input dimensions. Note the input dimension for index_add is two dimensional square arrays, while the input dimension for scatter_reduce is one dimensional.
Fig. 3 shows heat maps of $V_{c}$ as a function of $R$ and input dimension for the scatter_reduce and index_add operations with summation reductions. There is a clear trend of increased variability as a function of each of these parameters. In the case of larger input dimension, there are more opportunities for runtime non-determinism to occur. For small reduction ratios, when there are many source elements combined into a few output elements, there is less variation. We suspect this is due to the fact that our index tensor is random, meaning that elements being reduced will not be located locally in memory. For both operations we observe $V_{c}$ near one, meaning that most runs have a unique output. This is the worst case for reproducibility and error debugging.
@@
Figures 4 and 5 show $V_{\text{ermv}}$ and $V_{c}$ as a function of reduction ratio. For the scatter_reduce operation we use a one dimensional array with 2,000 elements, and for index_add we use a two dimensional square array with 100 $\times$ 100 elements. These array sizes are selected to be in an interesting regime of reduction ratio behavior as observed from the heat-maps. We notice a fairly constant $V_{c}\approx 0.1$ for scatter_reduce with both the sum and mean reductions. The $V_{c}$ for this operation at a reduction ratio of 1.0 has a value around 0.10 which is a significant jump. We notice similar behavior for $V_{\text{ermv}}$ for this operation. Note for the case where the output array has only one element, we recover the previous problem statement of computing the sum of an array, covered in Section III. For index_add, $V_{c}$ is increasing almost linearly with reduction ratio. The errors are inconsistently sized across reduction ratios, indicating different behavior at each reduction ratio. We also notice an approximately linear trend between $V_{\text{ermv}}$ and reduction ratio, with a particularly large error bar at $R=0.4$. The lack of trend for the errors on the variability requires further analysis.
TABLE V: Average kernel runtime for scatter_reduce and index_add kernels on the H100 and Groq using deterministic (D) and non-deterministic (ND) implementations.
Operation Implementation H100 $\left[\mu s\right]$ Groq $\left[\mu s\right]$
@@
We now investigate performance effects of deterministic and non-deterministic operations, performed on the H100 and LPU architectures. Up until now, the LPU architecture has been left out of the discussion of variability because its hardware operates deterministically [25]. In Table V we report the average kernel runtime for the scatter_reduce with input dimension 1,000 and reduction ratio $0.5$ and index_add with input dimension 1,000 $\times$ 1,000 and reduction ratio $0.5$ on both the H100 and LPU architectures. For these kernels we find the deterministic implementations on the H100 to be faster. This trend is not observed across all the operations in [24], however. For scatter_reduce and index_add the LPU architecture, which is by default deterministic, is faster than all GPU implementations. Runtimes for the complete set of functions in [24] are given in the GitHub [22].
V Effect of non-determinism on deep learning
Non-determinism in PyTorch kernels can adversely affect upstream tasks in model training and inference. To highlight this, we focus on training and inference for a Graph Sample and Aggregate (GraphSAGE) neural network.
@@ V-A GraphSAGE Convolution Network
Graph Neural Networks (GNNs) are extensively used for analyzing graph-structured data, such as social networks and molecular structures. GNNs operate on the principle of message passing, where each node in the graph aggregates information from its neighbors to update its own representation. The aggregation of information in software is often implemented via scatter and gather operations, which we have shown above have severe runtime variability. A layer in a GNN is defined by $h_{v}^{(k)}=\text{U}^{(k)}\left(h_{v}^{(k-1)},\text{ A}^{(k)}\left(\{h_{u}^{(k-1)}\mid u\in\mathcal{N}(v)\}\right)\right)$ where $h_{v}^{(k)}$ is the representation of node $v$ at layer $k$, $\mathcal{N}(v)$ denotes the set of neighbors of node $v$, and U and A are functions defining the update and aggregation steps, respectively. GraphSAGE (Graph Sample and AggregatE) is a popular graph convolution that leverages permutation invariant functions such as the sum and mean for the aggregation operation. Implementations of this often use the scatter_reduce and index_add operations discussed above, for example in the PyTorch Geometric [26] library. We now study the effect of non-determinism on training and inference for the GraphSAGE model.
@@ V-B Results
We trained a GNN with two SAGEConv layers on the CORA dataset. The Cora dataset is widely used for GNN research and benchmarks. CORA consists of 2,708 scientific publications classified into one of seven classes. The graph structure is created with 5,429 links representing citations between these publications. Each publication is described by a 1,433-dimensional feature vector. Training of the SAGEConv model is performed on this data set with a 10 epoch training run. The only source of non-determinism in our implementation of this DNN is the index_add operation. We studied the variability of model weights over 10 epochs and found that $V_{\text{ermv}}$ increased from 1.414(0.05) to 1.418(0.23) from epoch 1 to 10. This shows that as expected, non-determinism results in a compounding increase in the variability with more kernel calls. At the end of the training loop, all 1,000 models had a unique set of model weights ($V_{c}=1$). Despite this variability all models converge to similar loss values. Unlike stochastic training using random number generators for initialization, the resulting 1,000 models are completely non-reproducible, even for a single user on a single machine.
TABLE VI: $V_{\text{ermv}}$ and $V_{c}$ for different training and inference deterministic (D) and non-deterministic (ND) combinations on GPUs.
Training Inference $V_{\text{ermv}}\times 10^{-6}$ $V_{c}$

### v3

@@ III-B Other approaches: Parallel sums with OpenMP
OpenMP, a directive-based API for shared memory and accelerator parallel programming, supports data reduction to perform parallel calculations that are portable across architectures. The OpenMP programming model uses reduction scoping clauses to specify the regions of code where the reduction takes place. These regions can include parallel fork-join, tasks, sections, SIMD, and scope constructs [25], which can be executed on a host or a device such as a GPU. The OpenMP specification does not specify the location and ordering in which the values are combined, thus, bitwise determinism is not guaranteed. As a result, implementing a deterministic parallel reduction in OpenMP will require using constructs that can enforce ordering when reducing private and shared variables. An ordered construct can define a structured block within a loop, SIMD, or loop SIMD region. This construct enforces sequential execution for a specified region of code according to the loop iterations, while enabling parallel execution for code outside the region. When the thread handling the first iteration of the loop reaches the ordered construct, any other thread that encounters the ordered construct in its iteration will wait at the start of the ordered region until all ordered regions from previous iterations have been executed. To order the entire loop, an ordered clause can be used.
@@
Listing 2: An ordered parallel reduction implementation OpenMP
@@
Listing 3: An ordered parallel reduction implementation OpenMP using
@@
Listing 2 shows the use of a block-ordered directive on a reduction. A similar effect can be achieved while using an ordered clause. Listing 3 shows how to use an ordered clause inside a target region that executes on a device or GPU. When comparing the results of listing 2 with a reduction without the ordered directive, using gcc 12.2.0 on CPU, we get the results shown in Table III; the ordered version produces deterministic results.
TABLE III: Results of the normal and ordered reductions using OpenMP on CPU
@@
III-C Statistical properties of the variability of non-deterministic parallel sums using CUDA or HIP on different GPUs
It is often assumed that the variability due to FPNA can be described as Gaussian noise, but we could not find evidence supporting this in the literature. We therefore performed numerical experiments to estimate the probability density function (PDF) of $V_{s}(S)$. We generated a set of 100 arrays of one million FP64 elements each taken from the uniform distribution $U(0,10)$, applied SPA 10000 times to each array, and evaluated the variability $V_{s}$, using SPTR for the deterministic summation, returning one million values for $V_{s}$. $S_{\textrm{d}}$ thus refers here to SPTR while $S_{\textrm{nd}}$ refers to SPA. We then repeated the experiment, replacing the uniform distribution with a normal distribution of zero mean and standard deviation 1.
Figure 1: Probability density of the variability $V_{s}$ for sums of 1 M FP64 numbers sampled from normal and uniform distributions on the V100 GPU. Kernel parameters are $N_{t}=64$ and $N_{b}=7813$ for both SPTR and SPA.
The resulting PDF s are shown in Fig.1 for both distributions for the V100; for the Mi250X and GH200 GPUs the results can be found in the GitHub repository [24]. The means and standard deviations of $V_{s}$ are different between the GPU types, while the shapes are similar. Using Kullback–Leibler divergence criterion (KL) analysis, we find that all PDF s for SPA do converge towards a normal distribution whose mean and standard deviation depend on $x_{i}$ and the GPU family. However, when we repeated the experiments replacing SPA with AO on the NVIDIA GPUs, using a sample size of 500000 sums (results shown in Fig. 2 for the V100), the distribution is found not to be normal, showing that the assumption of Gaussian noise is invalid in general. The reasons for this distribution are unclear, as the NVIDIA runtime scheduler details are proprietary.
We also calculated the dependence $\text{Max}\left|V_{s}\right|$ as a function of the array size $n$ for the same sequences used for Fig. 1. The data can be fit with a power law of the form $\beta n^{\alpha}$. $\text{Max}\left|V_{s}\right|$ is proportional to $\sqrt{n}$ when $x_{i}\in U(0,10)$. The exponent is larger for $x_{i}\in\mathcal{N}(0,1)$, showing that the range of the numbers also plays a role.
Figure 2: PDF of the scalar variability $V_{s}$ for 1 M FP64 numbers sampled from the uniform distribution $U(0,10)$ when AO is used for the nondeterministic implementation, on V100.
III-D Performance comparisons
To illustrate the performance impact of these different strategies, we measured the time needed to compute 100 sums of 4194304 FP64 elements taken from the uniform distribution, using SPTR, TPRC, CU, AO and SPA. The only nondeterministic algorithms in this tests are AO and SPA. We compute the negative quantity $P_{s}=100(1-t_{i})/\text{min}(t_{i})$ where $\text{min}(t_{i})$ is the minimum of all timings $t_{i}$, to compare the performance of all implementations to the fastest one. The other parameters controlling the algorithm are the thread block size and the number of thread blocks. Details on the choice of these settings are provided in the associated repository [24].
The main results are summarized on Tab. IV. We find that AO is 2 orders of magnitude slower than the fastest implementations. The fastest sum implementation depends on the GPU version. The non-deterministic SPA seems to be favored on NVIDIA GPU, followed by SPTR and CU. The difference in timings between SPA and SPTR is less than 0.2% for the V100 GPU but can reach up to 7.8% on the GH200. CU suffers a 4.5% penalty on GH200. Most implementations on V100 are within 0.5% to 1% of each other. The performance penalty on GH200 is more spread than on V100. TPRC is the fastest implementation on the Mi250X GPU followed by CU.
TABLE IV: Timing and performance penalty of parallel sum implementations on different GPUs for 100 sums of 4194304 FP64 numbers and varying kernel parameters. Timings averaged over 10 consecutive runs; non-deterministic algorithms indicated in red. Standard deviations in parentheses.
GPU implementation ($N_{t}\times N_{b}$) time for 100 sums (in ms) $P_{s}$ (in %)
SPA $(512\times 128)$ $6.456(0.008)$ 00
SPTR $(512\times 128)$ $6.469(0.011)$ $-0.198538$
V100 TPRC $(512\times 128)$ $6.491(0.006)$ $-0.528162$
CU (unknown) $6.87697$ $-6.51331$
AO (fixed parameters) $872.004(0.022)$ $-28781.3$
SPA $(512\times 512)$ $3.019(0.022)$ 00
CU (unknown) $3.1553$ $-4.50533$
TPRC $(512\times 512)$ $3.226(0.030)$ $-6.87139$
GH200 SPTR $(512\times 512)$ $3.254(0.014)$ $-7.79916$
AO (fixed parameters) $738.687(0.037)$ $-24365.7$
TPRC $(512\times 256)$ $6.275(0.012)$ 00
CU (unknown) $6.378$ $-1.63577$
Mi250X SPA $(512\times 256)$ $6.394(0.032)$ $-1.90101$
SPTR $(256\times 512)$ $6.552(0.023)$ $-4.4171$
IV Non-determinism in PyTorch Functions
@@
Apart from the choice of kernel at runtime and atomic operations, there are several other ways in which PyTorch may induce variability. These include having an unset random number generator seed, having unset CUDA environment variables, uninitialized GPU memory and communication between devices. To focus exclusively on variability emergent from the first two sources, we a single GPU and remove those other sources of variability. We control whether or not non-deterministic kernels are available through the environment variable that PyTorch provides [26]. We note that this documentation may not necessarily be completely accurate across all systems and software versions, as we received a runtime error when trying to obtain a deterministic result for scatter_reduce, suggesting the potential difficulties high level users may experience when trying to control determinism within this deep software stack.
@@
We performed experiments in two ways, depending on whether or not a deterministic kernel exists. If there is a deterministic option, then the output multidimensional array $\mathbf{A}$ is fixed by that implementation. We compare it with the tensors $B_{i}$ with $i$ labelling $N$ runs of the non-deterministic implementation. If there is no non-deterministic kernel, we choose the first non-deterministic invocation as a reference, $\mathbf{A}=\mathbf{B}_{0}$. For each experiment we also present results for measurements of kernel runtime for non-deterministic implementations on the GPU, and deterministic implementations on the GPU and LPU architectures, to expose the performance costs of using deterministic operations. Runtime measurements are only for the execution time of the kernel, excluding data transfers to/from the device. For the GPU we make many measurements to report the average and standard deviation of the execution time. On the LPU architecture the runtime for the PyTorch function is reported as a fixed number since the cycle-by-cycle execution is determined ahead of time [27].
We performed a sweep over operation hyperparameters using 10000 runs on an H100 and observed only a handful of functions specified in the PyTorch documentation[26] to exhibit non-determinism. We list these operations in Table V alongside minimum and maximum $V_{\text{ermv}}$. We refer the reader to our repository [24] for in-depth details on the full hyperparameter sweep, where, for example, we consider kernel size, stride and padding when testing convolution operations.
@@
Our test results indicated that the main factor in ability to observe output variability in these kernels is the size of the input tensor, which gives more chances to observe the switching of order at runtime. For functions which perform reductions on an input tensor, we found that the ratio of the output dimension to input dimension is another key factor. We explore this further, restricting our analysis to scatter_reduce and index_add, which have such reductions.
@@
TABLE V: Max and min variability for non-deterministic pytorch operations over all hyperparameters tested.
Operation min$(V_{\text{ermv}})/10^{-7}$ max$(V_{\text{ermv}})/10^{-6}$
@@ IV-A Case studies of PyTorch kernels
The scatter_reduce function updates an output array $Y$ by applying a reduction operation on values from the input array $X$, according to indices specified in an index array $I$, for example, when reducing a one dimensional array with summation reduction, $Y[i]=\sum_{j}\left(\left\{X[j]\mid j=I[i]\right\}\right)$. This can be generalized to arrays with arbitrary dimensions and different reduction functions. The index_add operation updates the output array $Y$ by adding values from an input array $X$ according to the indices specified in an index array $I$. As an example, for two dimensional input and output arrays with a summation reduction over the first dimension, the elements of $Y$ are $Y[i,j]=Y[i,j]+X[k,j],\text{ for}\ i=I[k],\forall k$. This can similarly be generalized to arbitrary dimensions and different reductions other than addition. Both of these operations reduce the size of an input tensor, so we define the reduction ratio $R$ as the ratio of the output dimension size to source dimension size. When $R=1$, the arrays have the same size, and smaller values of $R$ correspond to larger reductions of the source array down to at most a singular value, along the axis of reduction. For the index tensors of both operations, we use random integers drawn from a uniform distribution to choose arbitrary values from the source tensor. The motivation for this is to mimic an arbitrary graph structure, where the reduction is happening over nodes that share an edge, although in this case we do not a priori expect any specific structure or hierarchy to be emergent.
Figure 3: Heatmaps of count variability ($V_{c}$) per run/iteration for 1,000 runs of the non-deterministic implementation of scatter_reduce  (left) and index_add  (right) for different reduction ratios and input dimensions. Note the input dimension for index_add is two dimensional square arrays, while the input dimension for scatter_reduce is one dimensional.
Fig. 3 shows heat maps of $V_{c}$ as a function of $R$ and input dimension for the scatter_reduce and index_add operations with summation reductions. There is a clear trend of increased variability as a function of each of these parameters. In the case of larger input dimension, there are more opportunities for runtime non-determinism to occur. For small reduction ratios there is less variation. We suspect this is due to the fact that our index tensor is random, meaning that elements being reduced will not be located locally in memory. For both operations we observe $V_{c}$ near one, meaning that most runs have a unique output. This is the worst case for reproducibility and error debugging.
@@
Figures 4 and 5 show $V_{\text{ermv}}$ and $V_{c}$ as a function of reduction ratio. For the scatter_reduce operation we use a one dimensional array with 2,000 elements, and for index_add we use a two dimensional square array with 100 $\times$ 100 elements. These array sizes are selected to be in an interesting regime of reduction ratio behavior as observed from the heat-maps. We notice a fairly constant $V_{c}$ between 0.005 and 0.01 for scatter_reduce with both the sum and mean reductions. The $V_{c}$ for this operation at a reduction ratio of 1.0 has a value around 0.10 (not plotted) which is a significant jump. We notice similar behavior for $V_{\text{ermv}}$ for this operation. Note for the case where the output array has only one element, we recover the previous problem statement of computing the sum of an array, covered in Section III. For index_add, $V_{c}$ is increasing almost linearly with reduction ratio. The errors are inconsistently sized across reduction ratios, indicating different behavior at each reduction ratio. We also notice an approximately linear trend between $V_{\text{ermv}}$ and reduction ratio, with a particularly large error bar at $R=0.4$. The lack of trend for the errors on the variability requires further analysis.
TABLE VI: Average kernel runtime for scatter_reduce and index_add kernels on the H100 and Groq using deterministic (D) and non-deterministic (ND) implementations. Standard deviations in parentheses.
Operation Implementation H100 $\left(\mu s\right)$ Groq $\left(\mu s\right)$
@@
We now investigate performance effects of deterministic and non-deterministic operations, performed on the H100 and LPU architectures. Up until now, the LPU architecture has been left out of the discussion of variability because its hardware operates deterministically [27]. In Table VI we report the average kernel runtime for the scatter_reduce with input dimension 1,000 and reduction ratio $0.5$ and index_add with input dimension 1,000 $\times$ 1,000 and reduction ratio $0.5$ on both the H100 and LPU architectures. For the latter we found the non-deterministic implementation on the H100 to be faster. This trend is not observed across all the operations in the documented list [26], however. For scatter_reduce and index_add the LPU architecture, which is by default deterministic, is faster than all GPU implementations. Runtimes for the complete set of functions in [26] are given in the supplemental repository [24].
V Effect of non-determinism on full deep learning workflows
Non-determinism in PyTorch kernels can adversely affect upstream tasks in model training and inference as found in real-world applications. To highlight this, we focus on training and inference for a Graph Sample and Aggregate (GraphSAGE) neural network.
@@ V-A GraphSAGE Convolution Network
Graph Neural Networks (GNNs) are extensively used for analyzing graph-structured data, such as social networks and molecular structures. GNNs operate on the principle of message passing, where each node in the graph aggregates information from its neighbors to update its own representation. The aggregation of information in software is often implemented via scatter and gather operations, which we have shown above have severe runtime variability. A layer in a GNN is defined by $h_{v}^{(k)}=\text{U}^{(k)}\left(h_{v}^{(k-1)},\text{ A}^{(k)}\left(\{h_{u}^{(k-1)}\mid u\in\mathcal{N}(v)\}\right)\right)$ where $h_{v}^{(k)}$ is the representation of node $v$ at layer $k$, $\mathcal{N}(v)$ denotes the set of neighbors of node $v$, and U and A are functions defining the update and aggregation steps, respectively. GraphSAGE is a popular graph convolution that uses functions such as the sum and mean for the aggregation. Implementations often use the scatter_reduce and index_add operations discussed above, for example in the PyTorch Geometric [28] library.
@@ V-B Results
We trained a GNN with two SAGEConv layers on the Cora dataset. This dataset is widely used for GNN research and benchmarks; it consists of 2,708 scientific publications classified into one of seven classes. The graph structure is created with 5,429 links representing citations between these publications. Each publication is described by a 1,433-dimensional feature vector. Training of the SAGEConv model is performed on this data set with a 10 epoch training run. The only source of non-determinism in our implementation of this DNN is the index_add operation. We studied the variability of model weights over 10 epochs and found that mean $V_{\text{ermv}}$ increased from from epoch 1 to 10, while the standard deviations also increased: 1.414(0.05) to 1.418(0.23). This may indicate that, as expected, non-determinism results in a compounding increase in the variability with more kernel calls. At the end of the training loop, all 1,000 models had a unique set of model weights: $V_{c}~{}1$. Despite this variability all models converge to similar loss values. Unlike stochastic training using random number generators for initialization, the resulting 1,000 models are completely non-reproducible, even for a single user on a single machine.
TABLE VII: $V_{\text{ermv}}$ and $V_{c}$ for different training and inference deterministic (D) and non-deterministic (ND) combinations on GPUs. Standard deviations in parentheses.
Training Inference $V_{\text{ermv}}/10^{-6}$ $V_{c}$